# FineFilter: A Fine-grained Noise Filtering Mechanism for Retrieval-Augmented Large Language Models

### v1

@@ 4.3 Adaptive Truncator
The goal of the adaptive truncator is to capture the minimal necessary clues based on the complexity of the question and the content of the retrieval documents, retaining sufficient clues to generate an accurate answer while further reducing reasoning overhead.
@@
To determine the optimal clues subset $\mathcal{S}^{t}$ for each query $q$, we perform data annotation based on the reranked answer clues $\mathcal{S}^{r}$ obtained from the previous re-ranking step. Given a query $q$ and its re-ranked clues $\mathcal{S}^{r}=\{s_{1}^{r},\dots,s_{n}^{r}\}$, the objective is to identify the smallest subset $\mathcal{S}^{t}$ such that the RAG system’s generation model $M$ can generate the correct answer $y$ based on $q$ and $\mathcal{S}^{t}$. We define $D_{k}=\{s_{1}^{r},\dots,s_{k}^{r}\}$, where $1\leq k\leq n$. The performance on each subset $D_{k}$ is evaluated by checking if the generation model’s output $M(q,D_{k})$ matches the ground truth $y$. The correctness condition is defined as:
@@
If the RAG system cannot generate a correct answer for any subset, then $\mathcal{S}^{t}=\emptyset$, indicating that no subset of the re-ranked sentences suffices to produce the correct answer. This method ensures that the minimal necessary context $\mathcal{S}^{t}$ is used, optimizing the balance between information relevance and computational efficiency.
During the model training stage, we fine-tune a LLM based on the data annotations as the adaptive truncator. The $Truncator$ is trained to predict the smallest index $K$ of $\mathcal{S}^{r}$ that needed to answer each query:
@@
During inference, given a new query $q$ and its reranked sentences $\mathcal{S}^{r}$, the $Truncator$ predicts the minimal index $K_{g}$ and truncates $\mathcal{S}^{r}$ to $\mathcal{S}^{t}=\{S^{r}_{1},\dots,S^{r}_{K_{g}}\}$. This ensures efficient utilization of computational resources while maintaining answer accuracy. Finally, the generation module of RAG concatenates the query $q$ with the filtered answer clues $\mathcal{S}^{t}$ as a prompt(see Appendix A.3) to reason the answer.
@@
Retrieval without Filtration
Top-1 document 36.81 69.21 5.17x 2.17 42.74 77.13 5.32x 3.11 25.54 60.09 4.83x 3.11
Top-5 documents 40.21 70.95 1.0x 1.69 48.32 80.16 1.0x 2.90 25.07 59.57 1.0x 1.82
@@
Retrieval with Filtration
RECOMP 37.12 69.43 11.97x 3.54 43.41 77.61 10.91x 3.25 24.59 59.26 12.95x 4.97
FILCO 32.43 64.78 17.43x 3.82 38.96 74.14 13.93x 3.47 20.12 56.03 11.77x 5.39
LongLLMLingua 36.96 69.25 4.56x 1.97 47.56 79.15 4.18x 3.04 24.31 58.93 4.45x 3.39
BottleNeck 39.72 70.14 14.32x 3.36 48.16 79.83 21.26x 4.32 25.64 60.23 13.21x 5.51
@@
Retrieval with Filtration
RECOMP 36.95 69.25 13.83x 3.25 43.39 77.51 10.91x 3.17 24.34 59.16 7.24x 4.35
FILCO 32.59 64.83 16.35 3.09 38.47 73.87 12.83x 3.31 21.34 56.91 13.00x 4.73
LongLLMLingua 37.45 69.67 4.09x 1.58 47.84 79.23 4.31x 3.01 24.05 58.75 4.22x 3.36
BottleNeck 39.48 70.05 12.53x 3.01 48.03 79.97 15.24x 4.28 25.47 59.97 11.06x 4.75
@@
Table 1: Experimental results on NQ, TriviaQA, and HotpotQA datasets. EM = exact match, F1 = F1 score, CR = compression ratio, TP = throughput (examples/second). We compare our FineFilter with Closed-book, Top-1, Top-5, and various filtering methods (RECOMP, FILCO, LongLLMLingua, BottleNeck) on two basement models LLaMA3-8B-Instruct and Mistral-7B-Instruct.
@@ 5.1 Experimental Setup
For the three open-domain QA datasets, we evaluate end-task performance using Exact Match (EM) and F1 score for the answer strings. EM measures exact correctness, while F1 evaluates answers that are close to but not necessarily exact, offering a more nuanced view of how well-predicted answers overlap with the correct ones. To assess the computational cost of downstream tasks, we introduce two metrics Cao et al. (2024); Hwang et al. (2024): compression ratio (CR) and inference throughput (TP) on a single A6000-48G GPU. The CR is defined as the ratio of the original context length to the compressed context length. TP refers to the number of examples the model can process or generate per second during inference.
@@
We use LLaMA3-8B-Instruct Dubey et al. (2024) and Mistral-7B-Instruct Jiang et al. (2023) as the backbone large language models. We fine-tune the two models with LORA Hu et al. (2021) as the clue extractor and adaptive truncator for 16 epochs on a single A6000-48G GPU. The initial learning rate is set to 5e-4, and the batch size is set to 4. We select the best model based on the performance of the validation set. For clue reranker, we implement Sentence-BERT Reimers and Gurevych (2020) using distilbert-base-uncased [^3] https://huggingface.co/distilbert/distilbert-base-uncased. In the final generation phase, we utilize the LLaMA2-7B Touvron et al. (2023) model for the three QA datasets.
@@ 5.2 Baselines
We select three types of baselines, including no filtration, extractive and abstractive filtration.
No Filtration
We evaluate the following three baselines: (i) Closed-book generation relying solely on parametric knowledge, (ii) Top-1 retrieval using the highest-ranked document, and (iii) Top-5 retrieval with direct concatenation of all retrieved documents.
@@
We choose RECOMP Xu et al. (2024) and LongLLMLingua Jiang et al. (2024). RECOMP employs a fine-tuned cross-encoder to identify salient sentences through dense retrieval. LongLLMLingua utilizes question-aware perplexity scoring with a dynamic programming algorithm to prune irrelevant tokens progressively in long contexts.
@@
We choose FILCO Wang et al. (2023) and BottleNeck Zhu et al. (2024). FILCO learns a context filtering model to dynamically identify key sentences and jointly optimizes with the generator for end-to-end content distillation. Bottleneck leverages reinforcement learning and information bottleneck theory to optimize filtering and generation.
@@ 5.3 Main Results
The comparison results on NQ, TrivialQA, and HotpotQA datasets are shown in Table 1. From the results, we can see that: RAG improves downstream task performance across three datasets. Using the top-5 documents generally outperforms using the top-1 document, indicating that incorporating more contextual information improves model performance. Noise Filtering are crucial for further improving performance. Across multiple datasets, applying filtering methods significantly reduces context length while maintaining performance close to that of the top-5 documents, effectively filtering out irrelevant information and enhancing accuracy. FineFilter outperforms baselines across multiple models and datasets. FineFilter consistently surpasses all filtration baselines across LLaMA3 and Mistral, achieving superior EM and F1 performance and compression efficiency, i.e., 6% and 37% improvement of EM and CR than BottleNeck on NQ dataset with LLaMA3. FineFilter performs remarkably better than Top-5 documents on complex multi-hop tasks. Compared to other datasets, FineFilter shows a larger improvement on the HotpotQA dataset, i.e., the EM improvement than Top-5 with LLaMA3 is 5.4% on HotpotQA, 4.8% on NQ and 1.0% on TriviaQA, highlighting its exceptional ability in handling complex multi-hop reasoning tasks.
Figure 4: An illustration of clue extraction model performance using LLaMA3-8B-Instruct on NQ, TriviaQA, and HotpotQA datasets. The x-axis represents the KNN-clustering threshold, with the model incorporating more context as the threshold increases.
Methods EM
FineFilter 42.17
@@
   w/o clue reranker 41.64
   w/o adaptive truncator 42.03
Table 2: The Exact Match of abalation study on NQ test set based on LLaMA3-8B-Instruct.
5.4 Analysis
5.4.1 Ablation Study
To explore the impact of different components on FineFilter, we use LLaMA3-8B-Instruct as the base LLM and introduce the following variants of FineFilter for ablation study:1) w/o clue extractor. LLaMA3-8B-Instruct, without fine-tuning, is used to directly extract answer clues; 2) w/o clue reranker. The original sentence ranking given by the fine-tuned extractor is used; 3) w/o adaptive truncator. No longer performs adaptive truncation on the re-ranked clues.
As shown in Table 5.3, removing the clue extractor and reranker leads to a significant drop in accuracy, highlighting the critical role of these components in the performance of FineFilter, with the clue extractor having a more pronounced impact on subsequent steps. In contrast, removing the truncator has a smaller impact on performance, as its contribution primarily lies in improving RAG reasoning efficiency.
Dataset Method EM
NQ Direct 39.41
Fine-tuned 41.43
@@ 5.4.1 Ablation Study
Fine-tuned 48.58
@@
Fine-tuned 26.12
Table 3: Performance comparison between Direct Extraction and Fine-tuned Extraction across three datasets based on LLaMA3-8B-Instruct.
5.4.2 Impact of Fine-tuning Clue Extractor
We analyze the impact of fine-tuning on the performance of the Clue Extraction module by comparing the following two approaches: 1) Direct Extraction. Using LLM without fine-tuning to extract clue sentences from retrieved documents based on extraction prompts (see Appendix A.1). 2) Fine-tuned Extraction. Fine-tune the extraction model to select the answer-containing sentences.
As shown in Table 3, the experimental results demonstrate that fine-tuned extraction outperforms direct extraction across all three datasets. Fine-tuning allows the model to select more relevant sentences by incorporating task-specific knowledge, leading to significant performance improvements.
5.4.3 Impact of KNN-Based Extraction
We evaluate the impact of the KNN-based extraction strategy by adjusting its threshold, which defines the distance of cosine similarity to the answer-containing sentences. When set to 0, the model selects sentences only containing the answer without KNN. As the threshold increases, the model gradually selects more semantically similar sentences to the answer, expanding the context and providing additional relevant information.
As shown in Figure 4, we compare the performance of the model at different threshold values. For simple questions such as NQ, the KNN extraction strategy does not improve performance, as answers can typically be obtained directly from sentences containing the answers. For more complex questions such as HotpotQA and TriviaQA, the KNN strategy initially improves performance at lower thresholds but declines at higher thresholds due to increased noise.
Methods EM F1
BM25 41.51 71.03
BGE-rerank 41.73 71.06
Sentence-BERT(Ours) 42.03 71.21
Dataset Method EM F1
NQ Random 41.23 71.01
Adaptive(Ours) 42.17 71.31
@@ 5.4.3 Impact of KNN-Based Extraction
Adaptive(Ours) 48.81 80.33
@@
Adaptive(Ours) 26.47 61.15
Table 5: Comparison of Random and our Adaptive Truncation methods across NQ, TriviaQA, and HotpotQA based on LLaMA3-8B-Instruct.
5.4.4 Impact of Different Reranker
To select a more effective re-ranking base model, we compare BM25 Robertson et al. (2009), BGE-rerank Xiao et al. (2024), and Sentence-BERTReimers and Gurevych (2020). As shown in Table 4, the results demonstrate that Sentence-BERT outperforms all baseline models in both EM and F1 scores, so we chose it as the re-ranking base model.
5.4.5 Adaptive vs. Random Truncation
To validate the effectiveness of our adaptive truncation strategy, we compare it with random truncation. As shown in Table 5, our adaptive truncation method outperforms random truncation in all metrics. This is because adaptive truncation dynamically selects context based on the results feedback from the downstream generator, retaining the most relevant and informative content to improve model performance.

### v4

@@ 4.3 Adaptive Truncator
The adaptive truncator aims to capture the minimal necessary clues based on the complexity of the question and the documents, ensuring sufficient clues for accurate answer generation.
@@
To determine the optimal clues subset $\mathcal{S}^{t}$ for each query $q$, we perform data annotation based on the reranked answer clues $\mathcal{S}^{r}$ obtained from the previous reranking step. Given a query $q$ and its reranked clues $\mathcal{S}^{r}=\{s_{1}^{r},\dots,s_{n}^{r}\}$, the objective is to identify the smallest subset $\mathcal{S}^{t}$ such that the RAG system’s generation model $M$ can generate the correct answer $y$ based on $q$ and $\mathcal{S}^{t}$. We define $D_{k}=\{s_{1}^{r},\dots,s_{k}^{r}\}$, where $1\leq k\leq n$. The performance on each subset $D_{k}$ is evaluated by checking if the generation model’s output $M(q,D_{k})$ matches the ground truth $y$. The correctness condition is defined as:
@@
If the RAG system cannot generate a correct answer from any subset, then $\mathcal{S}^{t}=\emptyset$, indicating no subset suffices. This method ensures the use of minimal necessary context $\mathcal{S}^{t}$.
During the model training stage, we fine-tune an LLM based on the data annotations. The $Truncator$ is trained to predict the smallest index $K$ of $\mathcal{S}^{r}$ that needed to answer the query:
@@
During inference, given a new query $q$ and its reranked sentences $\mathcal{S}^{r}$, the $Truncator$ predicts the minimal index $K_{g}$ and truncates $\mathcal{S}^{r}$ to $\mathcal{S}^{t}=\{S^{r}_{1},\dots,S^{r}_{K_{g}}\}$. Finally, the generation module of RAG concatenates the query $q$ with the filtered answer clues $\mathcal{S}^{t}$ as a prompt (see Appendix B.1) to reason the answer.
@@
Retrieval without Filtering
Top-1 document 36.81 69.21 5.17$\times$ 2.17 42.74 77.13 5.32$\times$ 3.11 25.54 60.09 4.83$\times$ 3.11
Top-5 documents 40.21 70.95 1.0$\times$ 1.69 48.32 80.16 1.0$\times$ 2.90 25.07 59.57 1.0$\times$ 1.82
@@
Retrieval with Filtering
RECOMPXu et al. (2024) 37.12 69.43 11.97$\times$ 3.54 43.41 77.61 10.91$\times$ 3.25 24.59 59.26 12.95$\times$ 4.97
LongLLMLinguaJiang et al. (2024) 36.96 69.25 4.56$\times$ 1.97 47.56 79.15 4.18$\times$ 3.04 24.31 58.93 4.45$\times$ 3.39
FILCO Wang et al. (2023) 32.43 64.78 17.43$\times$ 3.82 38.96 74.14 13.93$\times$ 3.47 20.12 56.03 11.77$\times$ 5.39
BottleNeck Zhu et al. (2024) 39.72 70.14 14.32$\times$ 3.36 48.16 79.83 $\boldsymbol{\times}$ 4.32 25.64 60.23 13.21$\times$ 5.51
@@
Retrieval with Filtering
RECOMPXu et al. (2024) 36.95 69.25 13.83$\times$ 3.25 43.39 77.51 10.91$\times$ 3.17 24.34 59.16 7.24$\times$ 4.35
LongLLMLinguaJiang et al. (2024) 37.45 69.67 4.09$\times$ 1.58 47.84 79.23 4.31$\times$ 3.01 24.05 58.75 4.22$\times$ 3.36
FILCO Wang et al. (2023) 32.59 64.83 16.35$\times$ 3.09 38.47 73.87 12.83$\times$ 3.31 21.34 56.91 13.00$\times$ 4.73
BottleNeck Zhu et al. (2024) 39.48 70.05 12.53$\times$ 3.01 48.03 79.97 15.24$\times$ 4.28 25.47 59.97 11.06$\times$ 4.75
@@
Table 1: Experimental results on NQ, TriviaQA, and HotpotQA using two base models: LLaMA3-8B-Instruct and Mistral-7B-Instruct.
@@ 5.1 Experimental Setup
We evaluate answer quality on three open-domain QA datasets using Exact Match (EM) for strict correctness and F1 score for partial overlap. To assess computational cost, we report compression ratio (CR) and inference throughput (TPCao et al. (2024); Hwang et al. (2024) on a single A6000-48G GPU. CR is defined as the ratio of original to compressed context length, and TP refers to the number of examples the generator can process or generate per second during inference.
@@
We use LLaMA3-8B-Instruct Dubey et al. (2024) and Mistral-7B-Instruct Jiang et al. (2023) as backbone LLMs. We fine-tune the two models with LORA Hu et al. (2021) as the clue extractor and adaptive truncator for 16 epochs on a single A6000-48G GPU. The initial learning rate is set to 5e-4, and the batch size is set to 4. We select the best model based on the performance of the validation set. For clue reranker, we implement Sentence-BERT Reimers and Gurevych (2020) using distilbert-base-uncased [^2] https://huggingface.co/distilbert/distilbert-base-uncased. In the final generation phase, we utilize the LLaMA2-7B Touvron et al. (2023) model. More details can be seen in Appendix A.
@@ 5.2 Baselines
We consider three baseline strategies:
Without Filtering
(i) Closed-book, which relies solely on the generator’s parametric knowledge; (ii) Top-1, which uses only the highest-ranked document for generation; (iii) Top-5, which concatenates the top five retrieved documents as input.
@@
(i) RECOMP Xu et al. (2024), which employs a fine-tuned cross-encoder to select salient sentences through dense retrieval; (ii) LongLLMLinguaJiang et al. (2024), which prunes irrelevant tokens in long contexts via a dynamic programming algorithm guided by question-aware perplexity scores.
@@
(i) FILCO Wang et al. (2023), which trains a context filtering model to dynamically select key sentences and jointly learns with the generator for end-to-end distillation; (ii) BottleNeck Zhu et al. (2024), which employs reinforcement learning and information bottleneck theory to improve both filtering and generation.
@@ 5.3 Main Results
The comparison results on three QA datasets are shown in Table 1. The results indicate the following: (i) RAG improves downstream task performance across all datasets. Using Top-5 documents generally outperforms Top-1, indicating that incorporating more contextual information improves model performance; (ii) Noise filtering is crucial for further improving performance. Across multiple datasets, filtering methods significantly reduce context length while preserving performance close to that of Top-5 documents, effectively removing irrelevant information and enhancing accuracy; (iii) FineFilter outperforms baselines across multiple models and datasets. FineFilter consistently outperforms all filtering baselines across LLaMA3 and Mistral, i.e., it improves EM by 6% and CR by 37% over BottleNeck on NQ with LLaMA3; (iv) FineFilter performs remarkably better than Top-5 documents on complex multi-hop tasks. FineFilter shows the largest improvement on HotpotQA, with a 5.4% EM gain over Top-5 using LLaMA3, followed by a 4.8% on NQ and 1.0% on TriviaQA, underscoring its superior performance in handling multi-hop QA.
LLaMA2-7B (NQ) Flan-T5-Large (HotpotQA)
Method EM F1 EM F1
FineFilter 42.17 71.31 23.79 58.44
@@
   w/o clue reranker 41.64 71.05 22.94 57.79
   w/o adaptive truncator 42.03 71.17 22.62 57.41
Table 2: Ablation study on NQ and HotpotQA test set. We use LLaMA2-7B and Flan-T5-Large as the generators, respectively.
Method EM $\uparrow$ F1 $\uparrow$ Latency (s.) $\downarrow$
LongLLMLingua 36.96 69.25 3.43
BottleNeck 39.72 70.14 3.63
RECOMP 37.12 69.43 0.97
FILCO 32.43 64.78 2.64
@@
Table 3: Latency Analysis. We report end-to-end inference latency and QA performance (EM, F1) on NQ test set using LLaMA3-8B-Instruct. Experiments are conducted on a single A6000 GPU.
Figure 4: An illustration of clue extraction performance using LLaMA3-8B-Instruct on NQ, TriviaQA, and HotpotQA. The x-axis shows the KNN threshold and higher values introduce more contextual sentences.
6 Analysis
In this section, we will conduct the ablation study, system latency evaluation, robustness analysis and generalization of FineFilter. Additional analysis can be seen in Appendix C and  D.
6.1 Ablation Study
To explore the impact of different components, we use LLaMA3-8B-Instruct as the base LLM and introduce the following variants of FineFilter for ablation study: 1) w/o clue extractor. directly uses LLaMA3-8B-Instruct without fine-tuning for clue extraction; 2) w/o clue reranker. kips reranking and retains the original sentence order; 3) w/o adaptive truncator. disables adaptive truncation of the reranked clues. As shown in Table 5.3, removing any single component leads to a performance drop in both EM and F1, confirming that all modules are essential to the overall effectiveness. Additional ablation results are presented in Appedix C.
Dataset Method EM F1
NQ Direct Extraction 39.41 69.98
Fine-tuned Extraction 41.43 71.01
@@ 6.1 Ablation Study
Fine-tuned Extraction 48.36 80.16
@@
Fine-tuned Extraction 26.01 60.73
Table 4: Performance comparison between Direct Extraction and Fine-tuned Extraction across three QA datasets using LLaMA3-8B-Instruct.
6.2 System Latency Evaluation
To assess the overall system efficiency, we measure the average total inference time (Latency, in seconds) required for all components in each method to process a single sample on the NQ test set. As shown in Table 3, RECOMP has the lowest latency but worse EM. FineFilter achieves the highest EM with only 0.52s more latency than the baselines, owing to its lightweight reranker and truncator that add minimal overhead, and the clue extractor with comparable complexity to existing methods. Thus, FineFilter offers a better trade-off between accuracy and efficiency.
6.3 Direct vs. Fine-tuned Extraction
We analyze the effect of fine-tuning on Clue Extraction by comparing two approaches: 1) Direct Extraction. Using LLM without fine-tuning to extract clue sentences from retrieved documents based on the given prompt (see Appendix B.2); 2) Fine-tuned Extraction. Using fine-tuned LLM to select answer-containing sentences. As shown in Table 4, Fine-tuned Extraction consistently outperforms Direct Extraction by leveraging task-specific knowledge to identify more relevant sentences.
6.4 Threshold of KNN-Based Extraction
We assess the KNN-based extraction by varying the threshold, which controls the cosine similarity to answer-containing sentences. A threshold of 0 selects only the answer sentences, while higher thresholds allow semantically similar sentences to expand the context with additional relevant information. As shown in Figure 4, we compare the model’s performance at different threshold values. For simple questions such as NQ, the KNN extraction strategy does not improve performance, as answers can typically be obtained directly from sentences containing the answers. For more complex questions such as HotpotQA and TriviaQA, the KNN strategy improves performance at lower thresholds but declines at higher thresholds due to increased noise. It is worth noting that KNN-based extraction plays a supplementary role. While the threshold setting impacts FineFilter’s performance, it does not change the experimental result that FineFilter outperforms the best noise filtering baseline.
6.5 Robustness Analysis
Cascading Errors Resilience Analysis
For high-quality retrieval scenarios, we conduct experiments on NQ test set with gold-standard answers to evaluate cascading errors. Recall-1 measures whether the extractor can retrieve the gold-standard answer. Hit-2@1 measures the probability that the reranker places the gold-standard answer as Top-1. Recall-3 measures whether the truncator can continue to retain the gold-standard answer. As shown in Table 5, FineFilter maintains a low error rate in the filtering components, indicating that cascading errors are kept within a controllable range.
Performance under Unreliable Retrieval
The truncator not only shortens context but also helps filter out unreliable contents. We conduct experiments on samples from HotpotQA test set without gold-standard answers to simulate unreliable retrieval. As shown in Table 6, FineFilter’s truncator effectively suppresses noise and prevents answer degradation. This is because FineFilter is trained to allow empty outputs, whereas baseline models perform poorly under unreliable retrieval conditions, as they must choose an output answer clue.
Dataset Recall-1 (%) Hit-2@1 (%) Recall-3 (%)
NQ 81.96 77.16 80.17
TriviaQA 91.94 83.25 85.28
HotpotQA 84.54 77.90 80.54
Table 5: Cascading errors analysis of extractor (Recall-1), reranker (Hit-2@1) and truncator (Recall-3) on three datasets using LLaMA3-8B-Instruct.
Method EM F1
RECOMP 15.51 51.33
FILCO 14.47 50.46
Ours (w/o Truncator) 15.63 51.42
Ours (w/ Truncator) 16.20 51.94
Table 6: Results under unreliable retrieval on HotpotQA test set using LLaMA3-8B-Instruct.
Method ROUGE-1 ROUGE-2 ROUGE-L
RECOMP 0.3051 0.1022 0.1668
FILCO 0.2743 0.0976 0.1428
BottleNeck 0.3708 0.1529 0.2158
LongLLMLingua 0.4082 0.1775 0.2369
Top-1 0.3532 0.1275 0.2103
Top-5 0.3769 0.1546 0.2234
@@ 6.5 Robustness Analysis
Table 7: Comparison of cross-task generalization ability across different baselines, evaluated using ROUGE-1, ROUGE-2, and ROUGE-L.
6.6 Cross-Task Generalization
To evaluate FineFilter’s generalization, we test it on 1,200 random samples from the Conversational Multi-Doc QA dataset  [^3] https://sites.google.com/view/wsdm24-docqa, after training on HotpotQA. As shown in Table 7, FineFilter maintains strong performance in this new dataset, demonstrating its generalization across diverse QA tasks.